# Victoria, BC, Canada

### v1

@@ 6.1 Presentation Antipatterns
  • Uninvited guest or party crasher: When one of the methods used is not clearly introduced in a paper introduction or methodology and makes an unexpected entrance in the later part of the paper, such as in the discussion or limitations sections. To fix this, the author should clearly mention all methods in the introduction and methodology sections of the paper, justifying why they were used. Even with emergent designs, as in the cases of Sam and Zara above, all methods used should be presented early in the paper with a mention that the design was emergent.
  • Smoke and mirrors: This antipattern can be spotted when a paper attempts to oversell a study as an MMMR when one of the approaches merely offers a token or anecdotal contribution to the research motivation or findings. This can be fixed by clarifying the presentation in the paper.
  • Limitation shirker: This antipattern occurs when there is a failure to discuss limitations from all methods used or from their integration in a study report. MMMR will have a much longer and more in depth limitation section than a single study paper, and the limitations reported should be congruent with the procedures used or mixed.
@@ 6.2 Study Design Antipatterns
  • Missing the mark: Misalignment of MMMR design with the research question/objective. This can be seen when one (or even both) of the methods do not make sense in light of the research question being asked. The fix for this should happen at the study design phase by asking if every method answers the core research questions being asked (whether planned up front or emergent). Note that sometimes the unit of observation may be changed but the unit of analysis remains the same. For example, Zara switched to looking at telemetry data rather than observing developers, but her unit of analysis remained the developers’ use of the tool for test generation.
@@
  • Selling your soul: This antipattern is related to “Missing the Mark” and may occur when a researcher employs an additional method to appeal to an anticipated reviewer who has preference for a particular methodology, but the use of the procedure does not contribute substantively to the research findings. This can be addressed by using the method one believes in or by ensuring that the researcher draws out meta-inferences from the multiple procedures used.
@@
  • Cargo cult research [^6] Cargo Cult can be a metaphor for superficial imitation of a process without basic understanding of its mechanism.: Using methods in which the research team lacks expertise but that they have observed being used to generate “exciting findings” in well cited papers. Collaborating with other researchers with broader expertise may help avoid this antipattern.
@@
  • Sample contamination: This antipattern may occur in an MMMR sequential design where the same participants are used in multiple, sequential methods without accounting for potential contamination from earlier method(s) to later ones (where participants have already seen an intervention or know about the research). However, in some cases the design itself may involve selecting participants (e.g., for interviews) from an earlier method (e.g., a survey) and the resulting “contamination” is intentional. Any possible contamination should be justified or reported, and potentially mentioned in limitations.
@@
  • Lost Opportunity: In an MMMR sequential design, when there is a failure to use findings from an earlier study during the development of an instrument for a study in a later phase of the research. This is an issue to be caught during the implementation of the study as it cannot be fixed through the presentation (alternatively, the studies can be written up as a two-phase study or two separate studies with two different contributions).
@@
  • Integration failure: We talked about failure to integrate data or insights as a quality attribute earlier in the paper. This antipattern occurs when there is poor integration of findings from all methods used in the study design. This can be fixed by planning how the results/findings will be integrated and making sure that the results/findings from both studies are integrated or merged in the results/findings section of the paper.
@@
  • Questionable Ethics: Finally, even when all of the above antipatterns are avoided, i.e., the MMMR study is applied and presented well, it may come across as having questionable aims, or as using questionable means to achieve the aims. The researcher should clearly state the motivation for the study and expected benefit to society or research to help justify the possible risks a study may involve.
@@ 7 Conclusions
In this paper, we provide an overview of how mixed and multi methods research designs can be used in software engineering, with the intention to inspire their use. Other resources, such as the ACM Sigsoft Empirical Standards sigsoft , as well as our earlier paper on selecting a research method in software engineering easterbrook_selecting_2008 , only lightly touch on the concepts and challenges of using mixed methods research in software engineering. As can be seen from Section 2, there is a long history of research behind the mixed method paradigm. Deciding what to include (and exclude) in this paper from this large body of research was challenging and we focused on the principles, characteristics and designs that we hope will support software engineering researchers in applying mixed methods to lead to impactful insights. Still, applying mixed methods is not a trivial exercise, as we try to demonstrate through the identification of many possible antipatterns.
@@
The “selecting empirical research methods” paper easterbrook_selecting_2008 notably elaborates on topics such as theoretical stance and the role of theory in quantitative, qualitative and mixed research. The main takeaways from that paper also apply here: a mixed or multi methods paper should have a clear research question that aligns with the research design; the procedures used should answer this question; the limitations of any one method or procedure can be used to motivate the choice of other methods, but the limitations of and threats associated with mixing should also be reported; and the role of theory and whether the research is inductive or deductive should drive the research design and shape the integrated findings reported.
@@
We also elaborate on four core principles of mixed and multi methods research that are: providing methodological rationale for the study design; ensuring the design leads to novel insights by the mixing of methods; assuring the rigorous use of individual methods and that they are mixed in a rigorous way; and finally, conducting the study in an ethical manner. These principles, at a high level, also apply to any research study, but we elaborate on how these principles can be considered specifically within the context of a mixed or multi methods research study.
@@
We further describe the core characteristics that shape a diverse landscape of mixed or multi methods research designs. These characteristics, which include the timing of procedures, the inductive or deductive nature of the research, the research design planning or emergence, how data is integrated, and the mixed or multi design, should be carefully considered up front and throughout the research process, and presented clearly in any research manuscripts. We do not provide a comprehensive description of all possible designs, but through several personas and research scenarios, we show how the principles and characteristics can be considered as tradeoffs and design choices are made. The integration process throughout the research design is what sets mixed methods apart.
@@
Our paper further suggests a number of practical considerations. For example, how to visualize the mixed or multi methods research design (as we show for the four personas), some potential pitfalls to try to avoid (articulated as research antipatterns), and which limitations to consider and present (guided in part by the principles). Mixed or multi methods, as mentioned, can be used to offset the limitations of a single method, and can be powerful for investigating socio-technical phenomena in software engineering, but they are challenging to use and often require broad expertise and substantial time commitments from the researchers involved.
@@
This paper is a first step in understanding how to use mixed methods in software engineering research and we hope that the SE research community will find this guide helpful. In closing, we suggest researchers adopt an open and creative stance to research design, and to share back with the community their insights gained not only on software engineering, but also on their research designs as they use mixed or multi methods.
@@
A research philosophy refers to beliefs, assumptions, values, and practices constituting a particular worldview or framework for understanding phenomena within a specific discipline or field of study (e.g., Positivism, Constructivism, Pragmatism, etc.). Paradigms provide a lens through which researchers interpret and make sense of reality, guiding their observations, theories, methods, and interpretations.
@@
Researchers specify the methods for analyzing the data collected during the study. These may involve statistical techniques, socio-technical grounded theory (STGT) for data analysis hoda_socio-technical_2021 , thematic analysis braun2012thematic , or other analytical approaches.
@@
Inductive and deductive reasoning are two complementary approaches used in research to develop theories, hypotheses, or generalizations. Inductive reasoning involves moving from specific observations or patterns to broader generalizations
Deductive reasoning involves moving from general principles or theories to predictions or hypotheses about specific instances. While inductive is commonly associated with quaL research methods, deductive is frequently associated with quaN research methods. However, this can vary, as exemplified in our scenarios (Section 5).
@@
Meta-inferences refer to making inferences or drawing conclusions about integrating the quaLand quaNmethods. When integrating quaLand quaNdata is a central element of the MMMR, how the researcher represents this integration and the interpretation of these findings becomes critical. During the MMMR discussion, the meta-inferences depicted during this process should be articulated to present how the use of an MMMR approach advanced a great understanding of the substantive topic compared to a mono-method approach.

### v2

@@ 6.1 Presentation Antipatterns
As mentioned, these antipatterns are concerns in how the research is presented or written. They are for the most part relatively easy to fix.
  • Uninvited guest or party crasher: This antipattern occurs when one of the methods used is not clearly introduced in a paper introduction or methodology and makes an unexpected entrance in the later part of the paper, such as in the discussion or limitations sections. To fix this, the author should clearly mention all methods in the introduction and methodology sections of the paper, justifying why they were used. Even with emergent designs, as in the case of Sam (Sec. 5.2), all methods used should be presented early in the paper with a mention that the design was emergent.
  • Smoke and mirrors: This antipattern can be spotted when a paper attempts to oversell a study as an MMR when one of the approaches merely offers a token or anecdotal contribution to the research motivation or findings. This can be fixed by clarifying the presentation in the paper. For example, in the case of the embedded design carried out by Zara (Sec. 5.4), if the treatment of the embedded part of the study (to find out how the tool impacted their confidence in testing) was shallow and the number of participants surveyed was very low, the paper would have oversold the contribution as a MMR study.
  • Limitation shirker: This antipattern occurs when there is a failure to discuss limitations from all methods used or from their integration in a study report. MMR will have a much longer and more in depth limitation section than a single study paper, and the limitations reported should cover all methods used and be congruent with the procedures mixed, as well as address how the data or analysis was integrated. In the case of Vicki (Sec. 5.3), failure to describe the limitations in how participants were observed (in this case, perhaps other forms of communication occurred outside Slack) would weaken the conclusions from integrating the insights from the observations with the commit behaviour.
@@ 6.2 Study Design Antipatterns
These antipatterns are ingrained in how the study was designed or conducted. They are hard to fix afterwards but thinking about them early on can help researchers avoid them.
@@
  • Missing the mark: Misalignment of MMR design with the research question or objective. This can be seen when one (or even both) of the methods do not make sense in light of the research question being asked. The fix for this should happen at the study design phase by asking if every method answers the core research questions being asked (whether planned up front or emergent). In the case of Zara, the research question considered not just the testing speed, but also student confidence. This could have been considered as two separate studies, but the second was done as an embedded study of convenience to broaden the insights collected.
@@
  • Selling your soul: This antipattern is related to “Missing the mark” and may occur when a researcher employs an additional method to appeal to a future reviewer who they believe has preference for a particular methodology, but the use of the procedure does not contribute substantively to the research findings. This can be addressed by using the method one believes in or by ensuring that the researcher draws out meta-inferences from the different procedures used. In the case of Ali (Sec. 5.1), the findings from the interviews may have been sufficient to stand on their own merit. But Ali may also have been influenced to do a survey to appease to a reviewer that may be focused on knowing if the findings generalized to more teams in the company.
@@
  • Cargo cult research [^8] Cargo Cult can be a metaphor for superficial imitation of a process without basic understanding of its mechanism.: Using methods in which the research team lacks expertise but that they have observed being used to generate “exciting findings” in well cited papers. Collaborating with other researchers with broader expertise may help avoid this antipattern. For example, in the case of Vicki adding a data mining method to the observation part of the method may require working with collaborators with mining expertise if Vicki lacks that expertise.
@@
  • Sample contamination: This antipattern may occur in an MMR sequential design where the same participants are used in multiple, sequential methods without accounting for potential contamination from earlier method(s) to later ones (where participants have already seen an intervention or know about the research). For example, in the case of Ali, Ali should be careful not to survey the participants already interviewed. However, in some cases the design itself may involve selecting participants (e.g., for interviews) from an earlier method (e.g., a survey) and the resulting “contamination” is intentional (for example in the case of a cross-over design). Any possible contamination should be justified or reported, and mentioned in limitations.
@@
  • Lost opportunity: In an MMR sequential design, when there is a failure to use findings from an earlier study during the development of an instrument for a study in a later phase of the research. This is an issue to be caught during the implementation of the study as it cannot be fixed through the presentation (alternatively, the studies can be written up as a two-phase study or two separate studies with two different contributions). In the case of Ali, not using insights from the interviews to ask pertinent questions in the survey about the security tool mandate could be a lost opportunity.
@@
  • Integration failure: We talked about failure to integrate data or insights as a quality attribute earlier in the paper. This antipattern occurs when there is poor integration of findings from all methods used in the study design. This can be fixed by planning how the results/findings will be integrated and making sure that the results/findings from both methods are integrated or merged in the results/findings section of the paper. In the case of Zara’s research, it may be difficult to see how to integrate the insights from the experiment and embedded survey. The insights of how the students became faster over time in the experiment, and were learning as gleaned from their self reports, can be integrated to emphasize the insight about learning using the automated testing tool and add credibility to the insights about effective testing from the dominant study. It may also be the case that the researcher integrates the findings from the mixed methods, but fails to adequately report them in the presentation (note, we could call this antipattern “Integration Presentation Failure”).
@@
  • Questionable ethics: Finally, even when all of the above antipatterns are avoided, i.e., the MMR study is applied and presented well, it may come across as having questionable aims, or as using questionable means to achieve the aims. The researcher should clearly state the motivation for the study and expected benefit to society or research to help justify the possible risks a study may involve. In the cases of all the fictional researchers above, we mentioned at least some of the ethical issues to be considered. All too often though we see papers were the respondents are not treated with this same care. We refer the reader to the points raised about conducting MMR ethically in Section 3.4.
7 Discussion
In this section, we address a number of issues and nuanced related to MMR. We start by expanding on how integration in MMR can be achieved and presented. We then discuss its synergies with other methods, including difference between mixed and multi methods research. Finally, we present a call for action.
7.1 Achieving Integration
Integration serves as a foundational element of MMR, as noted in Section 2 and depicted in Figure 1, and should be thoroughly detailed in the final reports of any MMR study. Creswell and Plano Clark endorse this view (poth_sage_2023, , Chapter 2) and explain that a comprehensive description of integration, considering both its intent and procedural aspects, enhances transparency by demonstrating how different methods were effectively employed. Below we present four types of integration in MMR, as recommended by Kuckartz and Rädiker (poth_sage_2023, , Chapter 22).
  • Sequential integration occurs when the findings from one method/study inform the application of another method/study. For instance, the results of an initial quaN method may guide the development of quaL data collection instruments or inform sampling strategies for the following phase. Sequential integration is seen in Ali and Sam’s scenarios (Sections 5.1 and 5.2).
  • Results-based integration involves combining, comparing, and relating findings only after the analyses of both quaL and quaN method have been completed. This type of integration focuses on synthesizing the outcomes of each method/study (quaL and quaN) to derive comprehensive insights. Results-based integration is seen in Vicki’s scenario (Sec. 5.3).
  • Data-based integration involves the simultaneous analysis and interrelation of quaL and quaN data during the analysis phase. This type of integration requires that both data types are associated with the same underlying source, ensuring that quaL and quaN data can be distinctly assigned and connected to each individual case. This alignment allows for a more cohesive and in-depth examination of the research subject. Data-based integration is seen in Zara’s scenario (Sec. 5.4).
  • Transformation-based integration occurs when quaL data is converted into quaN data, or vice versa, allowing for integration within a single “data world” (quaL or quaN). This type of integration enables researchers to analyze transformed data using the corresponding analytical techniques, enhancing the depth and breadth of the analysis. Transformation-based integration was not covered in our scenarios but examples can be found in (poth_sage_2023, , Chapter 22) and onwuegbuzie_routledge_2021 (dedicated book on how researchers conduct analysis in MMR).
Additional examples of integration can be found in Poth’s book poth_sage_2023 , including a dedicated section on “designing innovative integrations with technology” organized by Guetterman, who introduced the idea of joint displays during integration in MMR (see guetterman_integrating_2015 ; guetterman_visuals_2021 ). Di Penta and Tamburri Dipenta also propose the adoption of mind maps to gain an overview of the design to ensure there is alignment with the study’s quaN/quaL focus, and to aid in identifying gaps for further exploration, and to address threats to validity.
7.2 Synergies with Other Methods
Mixed methods research can easily be confused with multi methods research. The difference between mixed methods and multi methods lies in what types of data are used and how the data is used. Mixed methods requires the integration of quantitative (quaN) and qualitative (quaL) data. Multi methods can be applied with any combination including quaN-quaN or quaL-quaL. The principles we outline in Section 3 all apply to multi method research as well but we focus this paper on the use of the designs for quaN and quaL methods which can bring more emphasis on the socio-technical aspects in SE research.
Mixed methods are often used within or in combination with other research methods. For example, case studies runeson2012case , action research Staron20191 , socio-technical grounded theory hoda_socio-technical_2022 , ethnography sharp2016role , or experiment based studies wohlin2012experimentation can be conducted with a mixed methods research approach.
Finally, although the focus of this paper remains on primary studies, secondary research, such as systematic reviews of the literature (SLRs) Kitchenham2023SEGRESS ; PagePRISMA2020 , families of experiments Basili1999456 ; Santos2020566 , and rapid reviews Cartaxo2020 ; Bjarnason2023 that ground their application to methodological aspects and evidence synthesis using a mixed methods approach, can also benefit from our guidelines.
7.3 A Call to Action
In writing these guidelines, we do not simply promote MMR but emphasize using MMR ethically, as captured in the principle of ethical research. As a research community we have conducted large-scale surveys with developers and mined developer forums which in some cases has led to outrage in the developer community and introduction of new policies for unsolicited research calls for participation (see Section 3.4). As researchers, we should have empathy for the people we are studying, and indeed aim to help. They are real people. This is another reason why we include examples of fictional researchers, painting a picture of these researchers as human beings and the struggles they may face. Many of these struggles are not shared in the articles they write but perhaps could be so we can learn from each other. We recommend using MMR to make SE research more human focused – amplifying the human side of the quantitative data with the inclusion of qualitative data, both collected ethically.
As the lure of AI tools tempts the SE research community with promises of easing various research tasks and procedures, it is timely that we get into the habit of asking ourselves ‘why’, and not just be driven by the temptation of ‘why not’, as captured in our principle of methodological rationale. At the same time, as we begin to explore the risk and the opportunities of using LLMs for qualitative research bano2023LLM ; hoda2024Book , let us remind ourselves of the value of ‘deep work’, captured in the principle of procedural rigour. We strongly believe that using MMR studies conducted in interdisciplinary research settings can help address complex, socio-technical research questions and drive meaningful research, captured in our principle of novel integrated insights.
These guidelines for using MMR in SE are required now more than ever. We hope this will help present and future SE researchers use mixed methods research ethically, rationally, and rigorously to generate novel insights.
@@ 8 Conclusions
In this paper, we provide an overview of how mixed methods research designs can be used in software engineering, with the intention to inspire their use. Other resources, such as the ACM Sigsoft Empirical Standards sigsoft , as well as an earlier paper by Easterbrook et al. on selecting a research method in software engineering easterbrook_selecting_2008 , only lightly touch on the concepts and challenges of using mixed methods research in software engineering. As can be seen from Section 2, there is a long history of research behind the mixed method paradigm. Deciding what to include (and exclude) in this paper from this large body of research was challenging and we focused on the principles, characteristics and designs that we hope will support software engineering researchers in applying mixed methods to lead to impactful insights. Still, applying mixed methods is not a trivial exercise, as we try to demonstrate through the identification of many possible antipatterns. There is a lot more to read and learn about mixed methods, and the references we provide too can provide a starting point for researchers to learn more.
@@
The “selecting empirical research methods” paper easterbrook_selecting_2008 notably elaborates on topics such as theoretical stance and the role of theory in quantitative, qualitative and mixed research. The main takeaways from that paper also apply here: a mixed methods paper should have a clear research question that aligns with the research design; the procedures used should answer this question; the limitations of any one method or procedure can be used to motivate the choice of other methods, but the limitations of and threats associated with mixing should also be reported; and the role of theory and whether the research is inductive or deductive should drive the research design and shape the integrated findings reported.
@@
We also elaborate on four core principles of mixed methods research that are: providing methodological rationale for the study design; ensuring the design leads to novel integrated insights by the mixing of methods; assuring the rigorous use of individual methods and that they are mixed in a rigorous way; and finally, conducting the study in an ethical manner. These principles, at a high level, also apply to any research study (as alluded to above), but we elaborate on how these principles can be considered specifically within the context of a mixed methods research study.
@@
We further describe the core characteristics that shape a diverse landscape of mixed methods research designs. These characteristics, which include the timing of procedures, the inductive or deductive nature of the research, the research design planning or emergence, how data is integrated, and the mixed design, should be carefully considered up front and throughout the research process, and presented clearly in any research manuscripts. We do not provide a comprehensive description of all possible designs, but through several research scenarios, we show how the principles and characteristics can be considered as trade-offs and design choices to be made. The integration process throughout the research design is what sets mixed methods apart.
@@
Our paper further suggests a number of practical considerations. For example, how to visualize the mixed methods research design (as we show for the four scenarios), some potential pitfalls to try to avoid (articulated as research antipatterns), and which limitations to consider and present (guided in part by the principles and how integration should be achieved). Mixed methods, as mentioned, can be used to offset the limitations of a single method, and can be powerful for investigating socio-technical phenomena in software engineering, but they are challenging to use and often require broad expertise and substantial time commitments from the researchers involved.
@@
This paper is a first step in understanding how to use mixed methods in software engineering research and we hope that the SE research community will find this guide helpful. In closing, we suggest researchers adopt an open and creative stance to research design, and to share back with the community their insights gained not only on software engineering, but also on their research designs as they use mixed methods.
@@
A research philosophy refers to beliefs, assumptions, values, and practices constituting a particular worldview or framework for understanding phenomena within a specific discipline or field of study (e.g., Positivism, Constructivism, Pragmatism, etc.). Paradigms provide a lens through which researchers interpret and make sense of reality, guiding their observations, theories, methods, and interpretations easterbrook_selecting_2008 , (hoda2024Book, , Chapter 5)
@@
Researchers specify the methods for analyzing the data collected during the study. These may involve statistical techniques, socio-technical grounded theory (STGT) for data analysis hoda2024Book , thematic analysis braun2012thematic , or other analytical approaches.
@@
Inductive and deductive reasoning are two complementary approaches used in research to develop theories, hypotheses, or generalizations. Inductive reasoning involves moving from specific observations or patterns to broader generalizations. Deductive reasoning involves moving from general principles or theories to predictions or hypotheses about specific instances. While inductive is commonly associated with quaL research methods, deductive is frequently associated with quaN research methods. However, this can vary, as exemplified in our scenarios (Section 5).
@@
Meta-inferences refer to making inferences or drawing conclusions about integrating the quaL and quaN methods. When integrating quaLand quaNdata is a central element of the MMR, how the researcher represents this integration and the interpretation of these findings becomes critical. During the MMR discussion, the meta-inferences depicted during this process should be articulated to present how the use of an MMR approach advanced a great understanding of the substantive topic compared to a mono-method approach.
Multi-method study
A multi-method study may be planned or emergent and involves either two or more quaN procedures or two or more quaL procedures, but not a mix of both quaL and quaN. The main goal is typically to expand on the insights across procedures, and they may be carried out sequentially or in parallel. Distinct research methods can be used in a multi-method design.
Research method
A research method is a structured approach or systematic process used to investigate, collect, and analyze data to answer questions or solve problems related to software development and engineering. Common research methods include case studies, experiments, surveys, and ethnographic studies, each tailored to explore specific aspects of a research goal.